OpenStudy (anonymous):

Tom's surf shop estimates that advertising costs for one month will be given by the polynomial 6x^2+7x. He estimates that utilities costs for one month is given by 2x-5. Find and simplify an expression that would estimate Tom's total costs for one month.

OpenStudy (anonymous):

this is what i got; (6x^2+7x)+2x-5= 6x^2+7x+2x-5= 36x+9x-5= 45x+-5 can someone tell me if this is correct ?

OpenStudy (turingtest):

the \(x^2\) term cannot be combined with the x term also \(6x^2\) is not the same as \((6x)^2=36x^2\)

OpenStudy (turingtest):

so it is correct up to \[6x^2+7x+2x-5=...\]

OpenStudy (anonymous):

ok so now i can combine like terms, so 7x+2x=9x

OpenStudy (anonymous):

6x^2=36x

OpenStudy (anonymous):

oh 12x instead of 36x